OpenStudy (anonymous):

Find the distance between A(0,1) and B(-4,6) to the nearest tenth

OpenStudy (lieutenantgeneral):

the distance between A(0,1) and B(-4,6) calculates by the formula: 0 + 1 = 0 -4 + 6 = 2 2 - 0 = 2 The answer is 2

OpenStudy (anonymous):

a.4.5

OpenStudy (anonymous):

a.4.5 b.6.4 c.8.1 d.9.2

OpenStudy (anonymous):

@LieutenantGeneral what would the answer be ?

OpenStudy (anonymous):

@dan815 please help him so he will get off of here

OpenStudy (anonymous):

@I_Need_Help_With_Home can you help me ?